Circular No. 7156

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999by IN NGC 2841 A new apparent supernova has been reported independently by R. Ar-

bour, South Wonston, Hampshire, England, and by the Lick Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126). On behalf of LOSS, M. Papenkova, W. D. Li, and A. V. Filippenko, University of California at Berkeley, report that SN 1999by was found and confirmed on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 30.2 (mag ∼ 15.7) and May 1.2 UT (mag ∼ 15.0). Arbour’s unfiltered CCD discovery image (limiting mag ∼ 17.5), taken on Apr. 30.8749 with a 0.3-m f/6.3 Schmidt-Cassegrain reflector in the course of his supernova patrol, yields an estimate of the new star’s magnitude to be 15.1 and its offset to be 96′′ west and 86′′ north of the center of NGC 2841. Arbour forwards the following position obtained by M. Armstrong from an image obtained on Apr. 30.904: α = 9^{h}21^{m}52^{s}.07, δ = +51^{o}00′06′′.6 (equinox 2000.0). Papenkova et al. provide position end figures 52^{s}.18, 07′′.2, which is ∼ 100′′ west and 91′′ north of the galaxy nucleus; they add that there is a bright star (mag ∼ 11.1) located 11′′.7 west and 29′′.5 north of the supernova. A KAIT image of the same field on Apr. 25.2 (limiting mag ∼ 19.3) showed nothing at the position of the new star. Arbour found a prediscovery image taken on Apr. 28.8786 that shows SN 1999by at mag ∼ 16.0–16.5, though the supernova was not visible on images taken on Apr. 10, 14, 22, 23, or 27 (limiting mag ∼ 17.5). Papenkova et al. note that NGC 2841 is a well-known Sb galaxy and the host of three past supernovae (1912A, 1957A, 1972R), and they urge spectroscopic and photometric observations for this early and potentially bright supernova.

SUPERNOVA 1999bg IN IC 758 Additional CCD magnitude estimates, unfiltered unless noted other-

wise (cf. IAUC 7143): Apr. 19.030 UT, V = 15.85 (D. Hanžl, Brno, Czech Republic); 19.049, R = 15.43 (Hanžl); 29.531, 15.6 (S. Yoshida and K. Kadota, Ageo City, Japan); 29.808, R = 15.4 (Hanžl); 30.030, V = 15.85 ± 0.14 (Hanžl).

SUPERNOVA 1999bx IN NGC 6745 Note that the host galaxy of SN 1999bx is NGC 6745 = UGC 11391

(cf. IAUC 7154).

1999 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green